# OmniRouter: Budget and Performance Controllable Multi-LLM Routing

### v5

@@ Abstract
Large language models (LLMs) deliver superior performance but require substantial computational resources and operate with relatively low efficiency, while smaller models can efficiently handle simpler tasks with fewer resources. LLM routing is a crucial paradigm that dynamically selects the most suitable large language models from a pool of candidates to process diverse inputs, ensuring optimal resource utilization while maintaining response quality. Existing routing frameworks typically model this as a locally optimal decision-making problem, selecting the presumed best-fit LLM for each query individually, which overlook global budget constraints, resulting in ineffective resource allocation. To tackle this problem, we introduce OmniRouter, a fundamentally controllable routing framework for multi-LLM serving. Instead of making per-query greedy choices, OmniRouter models the routing task as a constrained optimization problem, assigning models that minimize total cost while ensuring the required performance level. Specifically, a hybrid retrieval-augmented predictor is designed to predict the capabilities and costs of LLMs. After obtaining the predicted cost and performance, we utilize a constrained optimizer for cost-optimal assignments that employs Lagrangian dual decomposition with adaptive multipliers. It iteratively converges toward the globally optimal query-model allocation, dynamically balancing latency minimization against quality thresholds while adhering to heterogeneous capacity constraints. Besides, we also introduce a new dataset, OmniRouteEval, which diverges from existing datasets that predominantly focus solely on either response quality or budget cost. Our dataset uniquely balances efficacy and efficiency considerations, containing sample-wise response quality and token usage measurements across 10 different LLMs, enabling comprehensive evaluation of both response quality and computational budget constraints. Experiments show that OmniRouter achieves up to 6.30% improvement in response accuracy while simultaneously reducing computational costs by at least 10.15% compared to competitive router baselines. The code and the dataset are available at https://github.com/agiresearch/OmniRouter.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities, powering a diverse range of applications from chatbots (Achiam et al., 2023; Team et al., 2023; Dubey et al., 2024; Guo et al., 2025) and code assistants (Hui et al., 2024; Wei et al., 2023; Nijkamp et al., 2023; Zhu et al., 2024) to computer-use agents. This success has spurred the widespread integration of LLMs into modern systems (Duan et al., ; Sun et al., 2024; Mei et al., 2024; Packer et al., 2023; Shi et al., 2025; Zheng et al., 2023b; Kwon et al., 2023). As LLM inference accelerates through hardware improvements, these LLM-integrated systems increasingly deploy not just one, but a set of LLMs with varying sizes, capabilities, and speeds as serving endpoints. This multi-LLM paradigm necessitates intelligent routing: the crucial task of directing incoming user queries to the most appropriate LLM instance to balance performance goals and resource efficiency.
@@
The design of effective LLM routers has become an active area of research. Recent works have proposed various routing strategies to optimize for cost (Ong et al., 2024; Somerstep et al., 2025; Anonymous, 2024), latency (Jin et al., 2023; Stojkovic et al., 2024; Zheng et al., 2024), and performance (Zhang et al., 2025; Yue et al., 2024; Chen et al., 2024; Zhuang et al., 2024). Most existing LLM routing frameworks predominantly treat routing as a sequence of independent, greedy decisions. For each incoming query, these routers select a model based on a local optimization criterion (e.g., lowest predicted latency, cheapest model predicted to succeed) without considering system-wide resource limitations or overall performance targets. These approaches fundamentally fail to achieve Pareto-efficient resource allocation across diverse query workloads distributed among multiple LLMs. When operating under budget constraints, these greedy routers cannot globally optimize the performance-resource tradeoff, leading to suboptimal overall system efficacy. For instance, expending computational resources on marginally improving responses to simple queries may leave insufficient capacity for complex queries where performance improvements would be more valuable, which can been in Figure 1. More critically, these localized decision frameworks cannot effectively enforce global constraints such as maintaining target response quality levels while adhering to strict resource budgets, resulting in either performance degradation or resource over-utilization when deployed at scale. In summary, these approaches confront a fundamental optimization dilemma: maintaining optimal overall performance while operating under computational resource constraints and stringent budgetary limitations.
@@
Figure 1: Comparison between traditional greedy routers and OmniRouter. Left: Greedy routers select models based on per-query optimization, leading to suboptimal allocations where Q1 (simple query) comes first and is assigned to the strong LLM, blocking Q2 (complex query) from accessing it. As a result, Q2 fails when assigned to the weak LLM. Right: OmniRouter employs constrained optimization to consider the global query distribution and model capabilities. It assigns the simple query to LLM1 (sufficient for the task) and reserves LLM2 for the complex query that requires more capability, thus maximizing overall success rate.
@@
Recognizing these limitations, we introduce OmniRouter, a controllable routing framework that fundamentally reimagines how query-model assignments should be determined. Unlike existing approaches, OmniRouter formalizes routing in multi-LLM systems as a constrained optimization problem (Zhang et al., 2020; Bertsekas, 2014; Homaifar et al., 1994) with global performance requirements and operational constraints and proposes a two-stage routing solution. Inspired by the success of retrieval-augmented generation (RAG) in reducing hallucination of LLMs, we design a retrieval-augmented predictor at the first stage to enhance the prediction-based router. At the second stage, we design the constrained optimizer which employs a dual gradient-based approach to navigate the solution space by adjusting model selections based on quality requirements and model concurrency. Our contributions are as follows:
$\bullet$  We introduce OmniRouteEval, a grounded dataset with query-model response correctness and token usage collected across 10 different models, enabling realistic evaluation of routing strategies.
@@ 2 Problem Formalization
Prior methods (Ong et al., 2024; Fu et al., 2024; Parkar et al., 2024) typically score or rank models for each query independently based on heuristics such as cost-effectiveness or response likelihood. While effective in simple settings, such formulations struggle to account for system-level constraints such as limited model concurrency, global quality targets, and budget bounds that are critical in real-world multi-LLM deployments. Different from existing works that formulate LLM routing as a ranking or selection problem, our approach models it as a constrained optimization problem.
@@
$\min_{x}\quad$ $\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}\quad s.t.\quad\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha\quad\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j\quad\sum_{j=1}^{M}x_{i,j}=1,\forall i$ (1)
It is a constrained optimization problem (Bertsekas, 2014; Homaifar et al., 1994) with global constraint $\alpha$ and local constraint $L$. We will elaborate how to solve this problem in the following section.
@@ 3 Methodology
As the capability $a_{i,j}$ and computational cost $c_{i,j}$ for each model-query pair are inherently uncertain at routing decision time. This uncertainty makes direct optimization of the routing variable $x_{i,j}$ particularly challenging. Joint optimization couples errors between prediction and allocation (Feng et al., 2023) and make the routing struggle with adaptability when query distributions shift or new models (Yu et al., 2022; Qin et al., 2023; Zhang et al., 2023). To address these challenges, we propose a two-stage approach that decouples the prediction of uncertain variables: $a_{i,j}$ and $c_{i,j}$ from the optimization of routing decisions $x_{i,j}$.
@@ 3.1 Retrieval-augmented Predictor
Inspired by retrieval-augmented generation (RAG) systems (Lewis et al., 2020; Borgeaud et al., 2022), which enhance model outputs by incorporating relevant retrieved information, we design a retrieval-augmented predictor, as illustrated in Figure 2, to integrate the generalization capabilities of trained models with the retrieval precision of historical query-model data.
We employ bert-base-uncased (Devlin, 2018) as our embedding encoder due to its representation capabilities and computational efficiency. This encoder processes both queries and LLM descriptions to generate embedding vectors $E_{q}$ and $E_{l}$ respectively, which serve as the foundation for both our training-based and retrieval-based prediction components. For the training-based component, we design a dual-head architecture built upon the embedding encoder. The first head focuses on model capability prediction, estimating a capability score $a_{i,j}^{pred}\in[0,1]$ for each query $i$ and model $j$ through:
@@
where $\sigma(\cdot)$ denotes the sigmoid activation function, $\mathbf{W}_{1}$ and $\mathbf{b}_{1}$ are learnable parameters. The second head performs sequence length classification by mapping the predicted output token length $l_{i,j}^{pred}$ into discrete buckets =$\mathcal{B}=\{B_{1},B_{2},...,B_{k}\}$, where $k=\lceil\frac{l_{max}}{bs}\rceil$, $l_{max}$ is the maximum sequence length, and $bs$ is the bucket size. The probability of length $l_{i,j}^{pred}$ is computed as:
@@
The bucketing strategy, a technique also employed in previous works (Jin et al., 2023; Zheng et al., 2024; Fu et al., 2024), transforms the challenging task of precise token-level prediction into a more tractable bucket-level approximation. During training, we optimize these dual objectives using mean squared error (MSE) for capability prediction and cross-entropy loss for length bucket classification.
@@
$l_{i,j}^{retrieve}=\frac{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})\cdot l_{m,j}}{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})},\quad a_{i,j}^{retrieve}=\frac{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})\cdot a_{m,j}}{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})}$ (4)
@@
$a_{i,j}=\gamma\cdot a_{i,j}^{pred}+(1-\gamma)\cdot a_{i,j}^{retrieve},\quad c_{i,j}=\delta\cdot tp_{j}(l_{i,j}^{pred})+(1-\delta)\cdot tp_{j}(l_{i,j}^{retrieve})$ (5)
@@ 3.2 Constrained Optimizer
At the second stage, we solve the $x_{i,j}$ with the predicted $a_{i,j}$ and $c_{i,j}$. Inspired by optimizers (Zhang et al., 2020; Bertsekas, 2014; Wang et al., 2023), we leverage the Lagrangian dual theory and introduce Lagrangian multipliers to convert the primal problem into its Lagrangian dual problem. We only provide the crucial steps in this section, and the complete induction can found in Appendix A. By introducing Lagrangian multipliers $\lambda_{1},\lambda_{2,j},\mu_{i}$, we get the Lagrangian relaxation function of the original problem as follows:
$L(x,\lambda_{1},\lambda_{2,j},\mu_{i})=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(-\frac{\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}}{N}+\alpha\right)$
$\quad\quad\quad\quad\quad\quad\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$ (6)
The dual function is formulated as Eq 7:
$g(\lambda_{1},\lambda_{2})=\min_{x_{i,j}}\mathcal{L}(x,\lambda_{1},\lambda_{2})=\min_{x_{i,j}}\{\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}(c_{i,j}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})+\lambda_{1}\alpha-\sum_{j=1}^{M}\lambda_{2,j}L_{j}\}$ (7)
and the dual problem is formulated as:
$\max_{\lambda_{1},\lambda_{2}}g(\lambda_{1},\lambda_{2})\quad\text{s.t.}\quad\lambda_{1}\geq 0,\lambda_{2,j}\geq 0,\forall j$ (8)
Then we can use gradient descent to update the partial derivative of $\lambda_{1}$ and $\lambda_{2}$, respectively, which is shown in Eq 9 and Eq 10.
$\lambda_{1}^{t+1}=\max(\lambda_{1}^{t}+\alpha_{1}(\frac{-\sum_{i=1}^{N}\sum_{j=1}^{M}x_{i,j}a_{i,j}}{N}+\alpha),0)$ (9)
$\lambda_{2,j}^{t+1}=\max(\lambda_{2,j}^{t}+\alpha_{2}(\sum_{i=1}^{N}x_{i,j}-L_{j}),0),\forall j\in[M]$ (10)
By updating $\lambda_{1}$ and $\lambda_{2}$, we can solve the partial sub-problem to obtain $x_{i,j}$ accordingly:
$x_{i,j}=\begin{cases}1&\text{if }j=j^{*}_{i}\\ 0&\text{otherwise}\end{cases}\quad j^{*}_{i}=\arg\min_{j\in[M]}(c_{ij}-\frac{\lambda_{1}a_{i,j}}{N}+\lambda_{2,j})$ (11)
An intuitive explanation of this process is: For $\lambda_{1}$: its partial derivative represents how much the quality constraint is violated. If the average quality is less than the threshold $\alpha$, the constraint is violated, and the gradient is positive, pushing $\lambda_{1}$ to increase, which in turn encourages the scheduler to select higher-quality models in the next iteration. For $\lambda_{2,j}$, its partial derivative shows the workload violation for each model j. If a model’s assigned queries exceed its capacity, the gradient is positive, increasing $\lambda_{2,j}$, which makes this overloaded model less attractive in subsequent iterations.
Proposition 1.
Given the Lagrangian dual problem with multipliers $\lambda_{1}\geq 0$ and $\lambda_{2,j}\geq 0$ for $j\in[M]$, the optimal assignment $x^{*}_{i,j}$ for the constrained optimization problem is uniquely determined by:
where $c_{i,j}$ and $a_{i,j}$ are the cost and capability scores predicted by the hybrid predictor, and $\lambda_{1}^{*}$ regulates the trade-off between efficiency and efficacy.
The optimal assignment rule shows that each query $i$ is routed to the model $j$ that minimizes the adjusted cost, where the adjustment term $\frac{\lambda_{1}^{*}a_{i,j}}{N}$ incentivizes selecting models with higher capability scores when $\lambda_{1}^{*}>0$.
Proposition 2.
Given the optimal dual variable $\lambda_{1}^{*}$ obtained from the gradient ascent update in Eq. (9), if $\lambda_{1}^{*}>0$, then the quality constraint is active and satisfied with equality:
$\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}^{*}=\alpha$ (13)
where $\alpha$ is the specified quality threshold in the constrained optimization problem.
This proposition ensures that the routing algorithm achieves exactly the required quality level $\alpha$ when the quality constraint is binding, preventing over-provisioning that would unnecessarily increase costs while maintaining the desired efficacy level. Detailed proof can be referred as Appendix A.
4 Evaluation
[…]
4.1 Datasets
Table 1: Comparison of OmniRouteEval with existing datasets used for LLM routing.
Dataset Token Usage Response Quality Cross-LLM Question Difficulty LMSYS-Chat (Zheng et al., 2023a) ShareGPT [^1] https://sharegpt.com/ RouterEval (Huang et al., 2025) OmniRouteEval
Since existing datasets for LLM routing can not satisfy the needs of routing under multiple constraints, we introduce OmniRouteEval, a dataset to facilitate the evaluation of LLM routing. The differences between OmniRouteEval and existing datasets for LLM routing are highlighted in Table 1. Within OmniRouteEval, we collect 2.7k questions sourced from established knowledge and mathematical reasoning datasets, including MMLU (Hendrycks et al., 2020), GPQA (Rein et al., 2023), MATH (Hendrycks et al., 2021), and GSM8K (Cobbe et al., 2021). We select 10 different models, including 5 relatively weak models, i.e., Qwen2.5 (7B-Instruct, 14B-Instruct, 32B-Instruct) (Yang et al., 2024), Gemma2 (9B-it, 27B-it) (Team et al., 2024) and 5 relatively strong models, i.e., Qwen2.5-72B-Instruct, GPT 4o-mini, GPT-4o (Achiam et al., 2023), Gemini-1.5-flash (Team et al., 2023), Claude-3.5-sonnet  [^2] https://claude.ai/. And we collect the response correctness and token usage of these models to answer these questions using Llama-3.1-70B-Instruct (Dubey et al., 2024) as the evaluation judge. The Llama-3.1-70B is excluded from the LLM candidate pool to reduce potential biases. Further details on the OmniRouteEval dataset construction and correctness assessment methodology are provided in Appendix B.
4.2 Setup
We conduct our experiments on an Ubuntu 22.04 machine equipped with 8 RTX A5000 GPUs to serve multiple LLMs using Ollama  [^3] https://ollama.com/. During the implementation, we set the maximum output length to 1024 tokens to prevent the influence of extremely long output sequences, set top $k$ as 16 during retrieval, and $\gamma$ and $\delta$ as 0.5 during aggregation. To calculate money costs, we refer to Litellm’s costmap for price reference. The detailed price map can be found in Appendix B. Our experiments are conducted on the continuous-batching setting (Kwon et al., 2023). The serving system processes requests by dynamically forming batches based on the incoming query traffic, providing a more realistic evaluation environment for routing frameworks. In our simulation, we model real-world traffic patterns by randomly adding $n\in\{1,2,3,4\}$ queries to the queue every 0.1 seconds and performing routing decisions at 1-second intervals. If not specifically mentioned, we set the two constraints $\alpha=0.75$ and the concurrent workload constraint $L=4$, which are applied uniformly across all LLMs. Controllability analysis of different values of $\alpha$ and $L$ is presented in Section 4.4. We evaluate OmniRouter against the following routing baselines:
Cost-oriented: We adapt S3 (Jin et al., 2023) and PO (Zheng et al., 2024), which are originally designed for latency optimization as cost-oriented baselines. Specifically, S3 employs DistilBERT (Sanh et al., 2019) and PO employs Vicuna-7B (Chiang et al., 2023) as output token length predictors, respectively.
Performance-oriented: For performance-focused routing, we implement EmbedLLM (Zhuang et al., 2024) and RouterDC (Chen et al., 2024). EmbedLLM leverages an encoder-decoder architecture to embed LLM representations, while RouterDC employs contrastive learning to model query-LLM relationships, both aiming to maximize response quality regardless of computational cost.
Cost-performance Coordinated: For baselines which balance both cost and performance objectives, we employ Hybrid-LLM (Anonymous, 2024), which constructs a probabilistic router, and CARROT (Somerstep et al., 2025), which implements dual predictors using Llama3-8B (Dubey et al., 2024) as specified in its original implementation.
4.3 RQ1: Routing Performance
The comparison of overall routing performance can be observed from Table 2. OmniRouter demonstrates substantial advantages over all baselines across both metrics. Given the theoretical performance bounds (lower bound is 57.41% when the worst LLM for each query is selected and upper bound is 90% when the best LLM for each query is selected), the improvement achieved by OmniRouter (1.3% to 6.48% absolute improvement) is particularly significant, which represents a meaningful advancement within this constrained optimization space. When compared to cost-oriented baselines (S3 and PO), OmniRouter achieves significantly higher success rates (5.74% and 6.48% improvements respectively) while maintaining comparable or even lower costs. Against performance-oriented approaches (EmbedLLM and RouterDC), our method achieves both higher success rates and substantially lower costs (approximately 41% cost reduction compared to RouterDC). Even when compared to cost-performance coordinated methods (Hybrid-LLM and CARROT), OmniRouter demonstrates superior performance in both dimensions.
Method Capability Acc. Length Bucket Acc. Exact Match ±$\pm 1$ S3 0.333 0.656 PO 0.325 0.683 EmbedLLM 0.732 RouterDC 0.761 Hybrid-LLM 0.708 CARROT 0.745 0.312 0.672 OmniRouter 0.813 0.452 0.806
Performance of Predictors. We also compare the effectiveness of OmniRouter’s predictor with the baselines we have mentioned. Table 3 demonstrates OmniRouter’s significant advantages in prediction accuracy. Our approach achieves 81.3% capability prediction accuracy, outperforming the strongest baseline (RouterDC) by 5.2%. For length prediction, OmniRouter delivers 45.2% exact match accuracy and 80.6% relaxed (±$\pm 1$) accuracy, surpassing all alternatives by over 13.0%.
4.4 RQ2: Controllability Analysis
(a) Cost vs. Performance Constraint
(b) Accuracy vs. Performance Constraint
(a) Cost vs. Concurrency Constraint
(b) Accuracy vs. Concurrency Constraint
4.5 RQ3: Ablation Studies
Predictor Variant Capability Acc ($\uparrow$) Bucket Acc ($\uparrow$) Accuracy ($\uparrow$) \$ Cost ($\downarrow$) Full 0.813 0.806 75.19% 0.0515 w/o Retrieval 0.751 0.724 71.85% 0.0698 w/o Training 0.728 0.682 73.33% 0.0642
Effects of Key Parameters within OmniRouter’s Predictors. We also investigate the sensitivity of OmniRouter’s routing effectiveness to key design parameters within its prediction components. However, due to the space limit, we move the results to Appendix C.
Broader Impact. By reducing the computational resources needed to achieve high-quality responses, OmniRouter contributes to more environmentally sustainable AI deployments with lower carbon footprints. While OmniRouter itself is designed to optimize resource allocation, the efficiency gains it provides might accelerate the proliferation of LLM deployments, potentially exacerbating privacy and security concerns associated with widespread AI adoption. We view these considerations not as arguments against developing more efficient routing systems, but rather as important areas for ongoing research and thoughtful implementation practices. Future work should explore how to incorporate fairness metrics directly into the constraint optimization framework.
5 Related Work
LLM Routing. To optimize performance and inference cost in multi-LLM serving systems, researchers have developed increasingly sophisticated routing approaches. RouteLLM (Ong et al., 2024) and CARROT (Somerstep et al., 2025) propose training small LLMs as routers to balance cost and performance. EmbedLLM (Zhuang et al., 2024) introduces a specialized encoder-decoder network for embedding LLM representations. HybridLLM (Anonymous, 2024) addresses data imbalance issues by proposing a probabilistic router to better represent different model capabilities across query types. C2MAB-V (Dai et al., 2024) treats routing as a contextual multi-armed bandit problem, enabling exploration-exploitation trade-offs in model selection. GraphRouter (Feng et al., 2024) leverages graph learning to jointly modeling the query-model, query-query, and model-model relationship for building routers. While these approaches have made substantial progress, they predominantly operate as greedy decision-makers, optimizing for individual queries without considering global system constraints or the implications for subsequent requests.
LLM Generation Length Prediction. Predicting LLM generation length is crucial for optimizing computational resources. Early attempts like Magnus (Cheng et al., 2024) employed random forest algorithms but achieved limited accuracy. Subsequent research has explored two main directions of prediction models: encoder-only models for classification (DynamoLLM (Stojkovic et al., 2024), S3 (Jin et al., 2023), TerriInfer (Hu et al., 2024), SSJF (Qiu et al., 2024b), and $\mu$3 (Qiu et al., 2024a)) and decoder-only models for generative prediction like Perception-only (PO) (Zheng et al., 2024). LTR (Fu et al., 2024) reformulated this as a ranking problem and utilized listwise ranking for predictor training.
6 Conclusion
Appendix
This appendix contains additional details for the paper: OmniRouter: Budget and Performance Controllable Multi-LLM Routing. The appendix is organized as follows: Section §A is about complete induction and proof of the constrained optimization. Section §B is about dataset details. Section §C is about additional experimental results. Section §D discusses more related works.
Appendix A Complete Induction and Proof of the Constrained Optimization
A.1 Complete Induction
The primal problem is as below:
@@ A.1 Complete Induction
s.t. $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha$ (15)
$\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j\in[M]$ (16)
$\sum_{j=1}^{M}x_{i,j}=1,\forall i\in[N].$ (17)
@@
$=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}+\alpha\right)$ (19)
$\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$ (20)

### v6

@@ Abstract
Large language models (LLMs) deliver superior performance but require substantial computational resources and operate with relatively low efficiency, while smaller models can efficiently handle simpler tasks with fewer resources. LLM routing is a crucial paradigm that dynamically selects the most suitable large language models from a pool of candidates to process diverse inputs, ensuring optimal resource utilization while maintaining response quality. Existing routing frameworks typically model this as a locally optimal decision-making problem, selecting the presumed best-fit LLM for each query individually, which overlooks global budget constraints, resulting in ineffective resource allocation. To tackle this problem, we introduce OmniRouter, a fundamentally controllable routing framework for multi-LLM serving. Instead of making per-query greedy choices, OmniRouter models the routing task as a constrained optimization problem, assigning models that minimize total cost while ensuring the required performance level. Specifically, a hybrid retrieval-augmented predictor is designed to predict the capabilities and costs of LLMs. After obtaining the predicted cost and performance, we utilize a constrained optimizer for cost-optimal assignments that employs Lagrangian dual decomposition with adaptive multipliers. It iteratively converges toward the globally optimal query-model allocation, dynamically balancing latency minimization against quality thresholds while adhering to heterogeneous capacity constraints. Experiments show that OmniRouter achieves up to 6.30% improvement in response accuracy while simultaneously reducing computational costs by at least 10.15% compared to competitive router baselines. The code and the dataset are available at https://github.com/dongyuanjushi/OmniRouter.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities, powering a diverse range of applications from chatbots [achiam2023gpt, team2023gemini, dubey2024llama, guo2025deepseek, hua2024trustagent, zhang2024ai, hua2024disentangling, iagent] and code assistants [hui2024qwen2, wei2023magicoder, nijkamp2023codegen2, zhu2024deepseek] to computer-use agents. This success has spurred the widespread integration of LLMs into modern systems [sun2024llumnix, mei2024aios, packer2023memgpt, shi2025from, zheng2023efficiently, kwon2023efficient, amem]. As LLM inference accelerates through hardware improvements, these LLM-integrated systems increasingly deploy not just one, but a set of LLMs with varying sizes, capabilities, and speeds as serving endpoints. This multi-LLM paradigm necessitates intelligent routing: the crucial task of directing incoming user queries to the most appropriate LLM instance to balance performance goals and resource efficiency. The design of effective LLM routers has become an active area of research. Recent works have proposed various routing strategies to optimize for cost [ong2024routellm, carrot2025, hybridllm2024], latency [jin2023s3, stojkovic2024dynamollm, panda2025adaptive, zheng2024response], and performance [zhang2025capability, jin2025two, chen2024routerdc, zhuang2024embedllm]. Most existing LLM routing frameworks predominantly treat routing as a sequence of independent, greedy decisions. For each incoming query, these routers select a model based on a local optimization criterion (e.g., lowest predicted latency, cheapest model predicted to succeed) without considering system-wide resource limitations or overall performance targets. These approaches fundamentally fail to achieve Pareto-efficient resource allocation across diverse query workloads distributed among multiple LLMs. When operating under budget constraints, these greedy routers cannot globally optimize the performance-resource tradeoff, leading to suboptimal overall system efficacy. For instance, expending computational resources on marginally improving responses to simple queries may leave insufficient capacity for complex queries where performance improvements would be more valuable, which can been in Figure 1. More critically, these localized decision frameworks cannot effectively enforce global constraints such as maintaining target response quality levels while adhering to strict resource budgets, resulting in either performance degradation or resource over-utilization when deployed at scale. In summary, these approaches confront a fundamental optimization dilemma: maintaining optimal overall performance while operating under computational resource constraints and stringent budgetary limitations.
@@
Figure 1: Comparison between traditional greedy routers and OmniRouter. Left: Greedy routers select models based on per-query optimization, leading to suboptimal allocations where Q1 (simple query) comes first and is assigned to the strong LLM, blocking Q2 (complex query) from accessing it. As a result, Q2 fails when assigned to the weak LLM. Right: OmniRouter employs constrained optimization to consider the global query distribution and model capabilities. It assigns the simple query to LLM1 (sufficient for the task) and reserves LLM2 for the complex query, thus maximizing overall success rate.
@@
Recognizing these limitations, we introduce OmniRouter, a controllable routing framework that fundamentally reimagines how query-model assignments should be determined. Unlike existing approaches, OmniRouter formalizes routing in multi-LLM systems as a constrained optimization problem [zhang2020solving, bertsekas2014constrained, homaifar1994constrained] with global performance requirements and operational constraints and proposes a two-stage routing solution. Inspired by the success of retrieval-augmented generation (RAG) in reducing hallucination of LLMs, we design a retrieval-augmented predictor at the first stage to enhance the prediction-based router. At the second stage, we design the constrained optimizer which employs a dual gradient-based approach to navigate the solution space by adjusting model selections based on quality requirements and model concurrency. Our contributions are as follows:
@@ 2 Problem Formalization
Prior methods [ong2024routellm, fu2024efficient, parkar2024selectllm] typically score or rank models for each query independently based on heuristics such as cost-effectiveness or response likelihood. While effective in simple settings, such formulations struggle to account for system-level constraints such as limited model concurrency, global quality targets, and budget bounds that are critical in real-world multi-LLM deployments. In contrast, our approach models it as a constrained optimization problem.
@@
$\min_{x}\quad$ $\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}\quad s.t.\quad\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha$ (1)
$\quad\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j\quad\sum_{j=1}^{M}x_{i,j}=1,\forall i$
It is a constrained optimization problem [bertsekas2014constrained, homaifar1994constrained] with global constraint $\alpha$ and local constraint $L$. We will elaborate how to solve this problem in the following section.
@@ 3 Methodology
As the capability $a_{i,j}$ and computational cost $c_{i,j}$ for each model-query pair are inherently uncertain at routing decision time. This uncertainty makes direct optimization of the routing variable $x_{i,j}$ particularly challenging. Joint optimization couples errors between prediction and allocation [feng2023towards] and make the routing struggle with adaptability when query distributions shift or new models [yu2022gilbo, qin2023towards, zhang2023s]. To address these challenges, we propose a two-stage approach that decouples the prediction of uncertain variables: $a_{i,j}$ and $c_{i,j}$ from the optimization of routing decisions $x_{i,j}$.
@@ 3.1 Retrieval-augmented Predictor
Inspired by retrieval-augmented generation (RAG) systems [lewis2020retrieval, borgeaud2022improving], which enhance model outputs by incorporating relevant retrieved information, we design a retrieval-augmented predictor, as illustrated in Figure 2, to integrate the generalization capabilities of trained models with the retrieval precision of historical query-model data.
We employ bert-base-uncased [devlin2018bert] as our embedding encoder due to its representation capabilities and computational efficiency. This encoder processes both queries and LLM descriptions to generate embedding vectors $E_{q}$ and $E_{l}$ respectively, which serve as the foundation for both our training-based and retrieval-based prediction components. For the training-based component, we design a dual-head architecture built upon the embedding encoder. The first head focuses on model capability prediction, estimating a capability score $a_{i,j}^{pred}\in[0,1]$ for each query $i$ and model $j$ through:
@@
where $\sigma(\cdot)$ denotes the sigmoid activation function, $\mathbf{W}_{1}$ and $\mathbf{b}_{1}$ are learnable parameters. The second head performs sequence length classification by mapping the predicted output token length $l_{i,j}^{pred}$ into discrete buckets $\mathcal{B}=\{B_{1},B_{2},...,B_{k}\}$, where $k=\lceil\frac{l_{max}}{bs}\rceil$, $l_{max}$ is the maximum sequence length, and $bs$ is the bucket size. The probability of length $l_{i,j}^{pred}$ is computed as:
@@
The bucketing strategy, a technique also employed in previous works [jin2023s3, zheng2024response, fu2024efficient], transforms the challenging task of precise token-level prediction into a more tractable bucket-level approximation. During training, we optimize these dual objectives using mean squared error (MSE) for capability prediction and cross-entropy loss for length bucket classification.
@@
$l_{i,j}^{retrieve}=\frac{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})\cdot l_{m,j}}{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})}$ (4)
$a_{i,j}^{retrieve}=\frac{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})\cdot a_{m,j}}{\sum_{q_{m}\in\mathcal{Q}_{k}(E_{q}^{i})}\text{sim}(E_{q}^{i},E_{q_{m}})}$ (5)
@@
$a_{i,j}=\gamma\cdot a_{i,j}^{pred}+(1-\gamma)\cdot a_{i,j}^{retrieve}$ (6)
$c_{i,j}=\delta\cdot tp_{j}(l_{i,j}^{pred})+(1-\delta)\cdot tp_{j}(l_{i,j}^{retrieve})$ (7)
@@ 3.2 Constrained Optimizer
At the second stage, we solve the $x_{i,j}$ with the predicted $a_{i,j}$ and $c_{i,j}$. Inspired by optimizers [zhang2020solving, bertsekas2014constrained, wang2023multi], we leverage the Lagrangian dual theory and introduce Lagrangian multipliers to convert the primal problem into its Lagrangian dual problem. By introducing Lagrangian multipliers $\lambda_{1},\lambda_{2,j},\mu_{i}$, we get the Lagrangian relaxation function of the original problem as follows:
@@
s.t. $\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}\geq\alpha$
$\sum_{i=1}^{N}x_{i,j}\leq L_{j},\forall j\in[M]$ $\sum_{j=1}^{M}x_{i,j}=1,\forall i\in[N].$
@@
$=\sum_{i=1}^{N}\sum_{j=1}^{M}c_{i,j}x_{i,j}+\lambda_{1}\left(-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{M}a_{i,j}x_{i,j}+\alpha\right)$
$\quad+\sum_{j=1}^{M}\lambda_{2,j}\left(\sum_{i=1}^{N}x_{i,j}-L_{j}\right)+\sum_{i=1}^{N}\mu_{i}\left(\sum_{j=1}^{M}x_{i,j}-1\right)$